OpenStudy (anonymous):

Create your own two-step equation with NO SOLUTION as the solution. Show the steps to solving the equation and explain in words why your work proves that there are NO SOLUTIONS

OpenStudy (anonymous):

Idk

OpenStudy (john_es):

This may be an example. Solve th equation, \[5x+3=5x-6\]Using the steps, \[5x-5x=-6-3\]Grouping numbers and variables in both sides \[0x=-9\]Summing, \[0=-9\] The last step shows you an imposible equality, so this shows us that this equation has no solutions. There no exists any x that satisfy the original equation given.

OpenStudy (anonymous):

I know but I want one that ends with fifteen. A two step equation